### v1

@@ Abstract
It was shown that pair luminosity of the newborn strange star with temperature of $10^{11}$ K may be as high as $L_{\pm}\simeq 10^{52}$ erg/s. The question remains: can a strange star maintain such a high surface temperature for a long time? To answer this question we studied thermal evolution of newborn strange star taking into account thermal conductivity of free quarks and neutrino emission by the URCA process. Our results show that extremely high luminosity due to the Schwinger process and insufficient thermal conductivity of quarks leads to development of steep temperature gradient at the surface of strange star. As a result, the temperature at the surface and hence its luminosity decreases as a power law, reaching $10^{40}$ erg/s already at 10 ms. This result holds even in the presence of neutrinosphere.
1 Introduction
@@
The strange quark matter (SQM) hypothesis (Bodmer, 1971; Witten, 1984) states that quark matter containing approximately equal numbers of up, down, and strange quarks is the true ground state of strongly interacting matter. From the SQM hypothesis a new class of astrophysical compact objects is predicted: quark stars or strange stars (Alcock et al., 1986; Weber, 2005) made up entirely of quark matter.
@@
Strange star can be formed as a result of phase transition in a neutron star (Berezhiani et al., 2003; Aguilera et al., 2004; Pagliara et al., 2013) or in a binary merger (Sedaghat et al., 2022). In any case, a newborn strange star is very hot. The temperature in the stellar interior can be as high as $10^{11}$ K, see Cheng & Dai (2001); Haensel et al. (1991); Drago et al. (2007). Powerful neutrino emission blows away the envelope of hadron matter, so the quark surface is bare and remains so as long as the temperature is higher than $\simeq 10^{7}$ K (Usov, 1997, 2001).
@@
Neutron combustion was shown to produce a huge energy release of up to $10^{53}$ ergs. Thus, strange stars are considered as possible progenitors of explosive phenomena such as supernovae and gamma-ray bursts (Olinto, 1987; Song, 2025; Berezhiani et al., 2003; Haensel & Zdunik, 2006) or fast radio bursts (Geng et al., 2021). At the same time, a newborn strange star can be a source of powerful pair emission with luminosity of about $10^{52}$ erg/s, as was shown in (Usov, 1998). The purpose of this paper is to consider this scenario in more detail.
@@
One of the unique features of strange stars is the existence of supercritical electric field in the thin surface layer of the star. The supercritical electric field exceeds the Schwinger limit for pair production $E_{c}=m_{e}^{2}c^{3}/\hbar e\simeq 1.3\times 10^{16}$ V/cm, where $m_{e}$ is the electron mass, $e$ is its charge, $c$ is the speed of light, and $\hbar$ is the reduced Planck constant. The region with supercritical electric field $E>E_{c}$ is called electrosphere. At low temperatures the electrosphere is static and does not produce electron-positron pairs due to the quantum degeneracy of electrons, i.e. the Pauli exclusion principle. On the contrary, hot electrosphere produces powerful pair wind (Usov, 1998; Prakapenia & Vereshchagin, 2024).
@@
It is crucial that in order to maintain high luminosity in pairs the surface temperature of bare strange star should remain high on a long timescale. For a gamma-ray bursts this is at least a few seconds. In the literature several models for the cooling of newborn strange stars were proposed, and it was argued that they can preserve the surface temperature about $10^{10}$ K for a few seconds (Berdermann et al., 2004; Li et al., 2021). Both these works use an integral approach with an average temperature assigned to a given volume of the strange star. At the same time, different approach was proposed by Page & Usov (2002) based on the heat transfer equation. The light curves were computed under the condition of neutrino transparency for a period of time starting from one second to $10^{10}$ seconds after strange star formation.
@@
The paper is organized as follows. In Section 2 we discuss an equation of state, hydrostatic state the heat transfer in hot quark star. In Section 3 we discuss electrostatic configuration and the physical conditions in the electrosphere. In Section 4 thermal evolution of newborn strange star is considered in both transparent and opaque regimes for neutrinos. Discussion and conclusion follow in the last Section.
2 Hydrostatic equilibrium configuration
@@
According to the bag model of non-interacting quark matter (Chodos et al., 1974) total pressure $P$ and total energy density $e$ are
@@
$P_{i}=\frac{g_{i}}{48\pi^{2}c^{3}\hbar^{3}}[\mu_{i}\left(\mu_{i}^{2}-m_{i}^{2}c^{4}\right)^{1/2}\left(2\mu_{i}^{2}-5m_{i}^{2}c^{4}\right)+$ (8)
$m_{i}^{4}c^{8}\ln\left(\frac{\mu_{i}+\sqrt{\mu_{i}^{2}-m_{i}^{2}c^{4}}}{m_{i}c^{2}}\right)],$
$\epsilon_{i}=\frac{g_{i}}{16\pi^{2}c^{3}\hbar^{3}}[\mu_{i}\left(\mu_{i}^{2}-m_{i}^{2}c^{4}\right)^{1/2}\left(2\mu_{i}^{2}-m_{i}^{2}c^{4}\right)$- (9)
$m_{i}^{4}c^{8}\ln\left(\frac{\mu_{i}+\sqrt{\mu_{i}^{2}-m_{i}^{2}c^{4}}}{m_{i}c^{2}}\right)].$
@@
Note that we consider non-interacting quark matter, where superconducting phase is absent. It was shown that quark pairing can increase $\mu_{e}$ which in turn increases the surface electric field. Another possibility to increase surface electric field without increasing $\mu_{e}$ is connected with the surface tension of quark matter.
Note, that the value of $\mu_{e}$ is almost unaffected by the temperature. Therefore we keep it constant during thermal evolution of the star.
3 Electrostatic configuration and Schwinger luminosity
In this Section, we describe the Schwinger process on the stellar surface, following Prakapenia & Vereshchagin (2024). As initial surface temperature is very high, the electric field binding electrons with quarks can produce pairs via the Schwinger process. Spatial distribution of electron density $n_{e}$ can be found from the Poisson equation, as described below. As it was pointed out by Prakapenia & Vereshchagin (2024), at finite temperature electrostatic configuration is not self-consistent, due to thermal evaporation of electrons. The corresponding stationary regime in hot dynamical electrosphere can be found by solving the Vlasov-Maxwell system.
@@ 3 Electrostatic configuration and Schwinger luminosity
which for high degeneracy (when $\eta\gg 1$) reduces to (Kettner et al., 1995; Usov et al., 2005)
@@
Poisson equation for quarks and electrons near the surface gives (Alcock et al., 1986; Kettner et al., 1995)
@@
To describe pair creation for non-vacuum initial state it is necessary to take into account Pauli blocking effect (Prakapenia & Vereshchagin, 2023; Brodin et al., 2023). We use the differential pair creation rate given by Gatoff et al. (1987) with additional Pauli blocking factor $(1-f_{e})$. The rate is then given by an integral over particle momentum $d^{3}p_{e}$ in the following form
@@
Figure 4: Normalized pair rate (16) (blue) with the corresponding normalized electric field (red) for $T=m_{e}$.
@@
The set of Vlasov-Maxwell equations for distribution functions $f_{\pm}$ and electric field $\tilde{E}$ in our case reduces to the Vlasov-Ampére system (Prakapenia & Vereshchagin, 2024)
@@
Figure 5: Stationary solution of Vlasov-Maxwell equations for electrosphere for $T=2m_{e}$ with dependence on distance of electric fiedl (top), particle density (middle) and luminosity (bottom). Approximate electrostatic solution given by eq. ((15)) is shown in black. Colors correspond to: electric field (red), electrons (blue) positrons (orange), andelectrostatic solution (black).
@@
4 Heat transfer and neutrino losses
@@
Now we are going to consider heat transfer in the quark star accounting for temperature evolution on much large timescales than the ones studied above. First we treat neutrino transparent case, following Page & Usov (2002).
Specific heat capacity $c_{v}$ for degenerate quarks is (Heiselberg & Pethick, 1993)
@@
Heat conductivity $\kappa$ for unpaired degenerate quarks is (Heiselberg & Pethick, 1993)
$\kappa=\frac{\pi^{4}k_{B}}{6\zeta(3)ch^{2}\alpha_{s}}\mu^{2},$ (21)
where $\alpha_{s}=0.3$ is QCD coupling constant.
@@
4.1 Neutrino transparent regime
@@
We solved numerically equation (22) for the stellar configuration with the mass $M=1.4M_{\odot}$. The temperature evolution of the star is shown in figure 6 together with luminosity evolution. After the time moment $t>1$ second, our curves demonstrate the same shape as in fig. 1 in (Page & Usov, 2002). Luminosity also decreases faster after $t>10^{5}$ seconds.
@@
The most important result is that the temperature evolution on the surface is independent on interior of the star. Surface temperature decreases as a power law up to few $10^{9}$ K at $t\simeq 10^{-2}$ seconds where it stabilizes. The corresponding luminosity is about $10^{40}$ erg/s. The temperature gradient disappears only at $10^{5}$ seconds, when the pair luminosity starts to decrease rapidly.
The total energy emitted in neutrinos is about $10^{53}$ ergs, while the total energy emitted in pairs is only $4\times 10^{41}$ ergs.
@@
4.2 Neutrino opaque regime
It is believed that neutrinos cannot escape freely immediately after strange star formation, until internal temperature decreases below $10^{10}$ K (Page & Usov, 2002). Following that idea, the integral models for neutrinosphere was developed by Berdermann et al. (2004); Li et al. (2021), where quark star is divided into two parts having different temperatures, separated from each other by a neutrinosphere with radius $R_{\nu}<R$. Neutrinosphere emits blackbody neutrino spectrum from its surface. The results by Berdermann et al. (2004) imply that neutrino opacity delays the process of cooling by an order of magnitude when compared to the transparent case. The results in (Li et al., 2021) imply that high luminosity in pairs up to $10^{50}$ erg/s can be sustained for long time, up to $10$ seconds. In this Section we compare the integral model developed in (Li et al., 2021) with the differential models adopted in our work.
Figure 7: Surface temperature evolution for neutrino transparent shell of the star with selected positions of neutrinosphere: $R-R_{\nu}=0.01$ cm (red), $R-R_{\nu}=0.1$ cm (green), $R-R_{\nu}=1$ cm (blue), $R-R_{\nu}=10$ cm (orange), $R-R_{\nu}=100$ cm (brown). Black curve represents blue curve from figure 6 (top) for convenience.
@@ 4.2 Neutrino opaque regime
Consider an outer shell of the star and fix the temperature at its inner boundary to be equal to the initial temperature $T(t,r=R_{\nu})=T(t=0,r)=T_{0}$. We solve numerically the heat transfer equation for different thickness of the shell $R-R_{\nu}=\{0.01,0.1,1,10,100\}$ cm. The results for the surface temperature are presented in figure 7. It is clear that the surface temperature decreases very fast in all cases. When the thickness increases up to $R-R_{\nu}=100$ cm cooling occurs in the the same was as for neutrino transparent star, that is the surface temperature is below $10^{10}$ K at the time moment $t\simeq 10^{-7}$ seconds. Finally, in figure 8 we show temperature distribution $T(r)$ and heat flux distribution $F_{r}(r)$ (for convenience we plot the value $4\pi R^{2}F_{r}$) inside the shell with thickness about 1 cm at different instants of time. Clearly, initially very steep temperature gradient emerges instantly and remains localized at the surface, due to weak thermal conductivity of quarks.
@@
In realistic model neutrinosphere moves toward the center of the star and at the same time loses its temperature. It means that our model overestimates the surface temperature. Thus, our results imply that the surface temperature of a newborn strange star made of noninteracting quarks decreases below the value $10^{10}$ K on the timescales $\Delta t\ll 1$ second, and the reason for that is the very high luminosity due to the Schwinger process and weak thermal conductivity of free quarks.
5 Discussion and conclusions

### v2

@@ Abstract
It was shown that pair luminosity of the newborn strange star with temperature of $10^{11}$ K may be as high as $L_{\pm}\simeq 10^{52}$ erg/s. The question remains: can a strange star maintain such a high surface temperature for a long time? To answer this question we studied thermal evolution of newborn strange star made of unpaired quarks taking into account its thermal conductivity and neutrino emission by the URCA processes $d\rightarrow u+e+\bar{\nu}_{e}$ and $u+e\rightarrow d+\nu_{e}$. Our results show that extremely high luminosity due to the Schwinger process and insufficient thermal conductivity of quarks leads to development of steep temperature gradient at the surface of strange star. As a result, the temperature at the surface and hence its luminosity decreases, reaching $10^{43}$ erg/s already at $10^{2}$ seconds. This result holds even in the presence of neutrinosphere.
I Introduction
@@
The strange quark matter (SQM) hypothesis 1971PhRvD…4.1601B ; 1984PhRvD..30..272W states that quark matter containing approximately equal numbers of up, down, and strange quarks is the true ground state of strongly interacting matter. From the SQM hypothesis a new class of astrophysical compact objects is predicted: quark stars or strange stars 1986ApJ…310..261A ; 2005PrPNP..54..193W made up entirely of quark matter.
@@
Strange star can be formed as a result of phase transition in a neutron star 2003ApJ…586.1250B ; 2004A&A…416..991A ; 2013PhRvD..87j3007P or in a binary merger 2022PhLB..83337388S . In any case, a newborn strange star is very hot. The temperature in the stellar interior can be as high as $10^{11}$ K, see chengDai2001 ; 1991ApJ…375..209H ; 2007ApJ…659.1519D . Powerful neutrino emission blows away the envelope of hadron matter, so the quark surface is bare and remains so as long as the temperature is higher than $\simeq 10^{7}$ K Usov_1997 ; Usov_2001 .
@@
Neutron combustion was shown to produce a huge energy release of up to $10^{53}$ ergs. Thus, strange stars are considered as possible progenitors of explosive phenomena such as supernovae and gamma-ray bursts OLINTO198771 ; PhysRevD.111.063040 ; 2003ApJ…586.1250B ; Haensel:2006kkg or fast radio bursts 2021Innov…200152G . At the same time, a newborn strange star can be a source of powerful pair emission with luminosity of about $10^{52}$ erg/s, as was shown in 1998PhRvL..80..230U . The purpose of this paper is to consider this scenario in more detail.
@@
One of the unique features of strange stars is the existence of supercritical electric field in the thin (few hundred fm) layer on its surface 1986ApJ…310..261A . This layer is called electrosphere 2005ApJ…620..915U . The strength of electric field $E$ there exceeds the Schwinger limit for pair production $E_{c}=m_{e}^{2}c^{3}/\hbar e\simeq 1.3\times 10^{16}$ V/cm, where $m_{e}$ is the electron mass, $e$ is its charge, $c$ is the speed of light, and $\hbar$ is the reduced Planck constant. The field strength depends on the surface density gradient 2010JPhG…37g5201M and may reach values of $5\times 10^{17}$ V/cm 1986ApJ…310..261A ; 1995PhRvD..51.1440K . At low temperatures Schwigner mechanism of pair production does not operate due to Pauli blocking of electronic states. Usov 1998PhRvL..80..230U have shown that electrosphere of newly formed hot quark star produces powerful electron-positron wind. Recently we reconsidered 2024ApJ…963..149P the mechanism of pair production in electrosphere of quark stars.
@@
It is crucial that in order to maintain high luminosity in pairs the surface temperature of bare strange star should remain high on a long timescale. For a gamma-ray bursts this is at least a few seconds. In the literature several models for the cooling of newborn strange stars were proposed, and it was argued that they can preserve the surface temperature about $10^{10}$ K for a few seconds 2004csqn.conf..399B ; 2021ApJ…922..214L . Both these works use an integral approach with an average temperature assigned to a given volume of the strange star. At the same time, different approach was proposed in 2002PhRvL..89m1101P based on the heat transfer equation. The light curves were computed under the condition of neutrino transparency for a period of time starting from one second to $10^{10}$ seconds after strange star formation.
@@
The paper is organized as follows. In Section II we discuss an equation of state, hydrostatic state the heat transfer in hot quark star. In Section III we discuss electrostatic configuration and the physical conditions in the electrosphere. In Section IV thermal evolution of newborn strange star is considered in both transparent and opaque regimes for neutrinos. Discussion and conclusion follow in the last Section.
II Hydrostatic equilibrium configuration
@@
According to the bag model of non-interacting quark matter ( i.e. neglecting QCD corrections and color superconductivity FahriJaffe1984 ; 1995PhRvD..51.1440K ) total pressure $P$ and total energy density $\epsilon$ are
@@
$P_{i}=\frac{g_{i}}{48\pi^{2}c^{3}\hbar^{3}}\left[\mu_{i}\left(\mu_{i}^{2}-m_{i}^{2}c^{4}\right)^{1/2}\left(2\mu_{i}^{2}-5m_{i}^{2}c^{4}\right)+\right$. (8)
$\left.m_{i}^{4}c^{8}\ln\left(\frac{\mu_{i}+\sqrt{\mu_{i}^{2}-m_{i}^{2}c^{4}}}{m_{i}c^{2}}\right)\right],$
[…]
$\left.m_{i}^{4}c^{8}\ln\left(\frac{\mu_{i}+\sqrt{\mu_{i}^{2}-m_{i}^{2}c^{4}}}{m_{i}c^{2}}\right)\right].$
@@
In this paper we consider SQM without color superconductivity and focus on the surface properties of the star. We calculated average quark chemical potential on the surface $\mu=(\mu_{u}+\mu_{d}+\mu_{s})/3=293.8$ MeV. This value determines the surface chemical potential of electrons and as a result the maximal surface electric field.
Including QCD corrections in the equation of state can only slightly increase the value of $\mu$ Alford_2005 . Another possibility is connected with the surface tension of quark matter. As shown in 2005ApJ…620..915U the surface electric field can be 10 times larger than the surface electric field calculated for SQM in the unpaired phase without the surface effects.
We can omit the last effect, because as we show in section III the Schwinger rate contains the Pauli blocking. Therefore, increasing the chemical potential will increase the electric field and enhance Pauli blocking effect. Thus, the resulting Schwinger rate remains almost unchanged.
As we show in section IV the key ingredient is thermal conductivity of quark matter, because it affects the temperature gradient on the surface induced by huge pair luminosity.
Finally, the value of $\mu$ is almost unaffected by the temperature 1995PhRvD..51.1440K . So, we keep it constant during thermal evolution of the surface.
III Electrostatic configuration and Schwinger luminosity
In this Section, we describe the Schwinger process on the stellar surface, following 2024ApJ…963..149P . As initial surface temperature is very high, the electric field binding electrons with quarks can produce pairs via the Schwinger process. Spatial distribution of electron density $n_{e}$ can be found from the Poisson equation, as described below. As it was pointed out in 2024ApJ…963..149P , at finite temperature electrostatic configuration is not self-consistent, due to thermal evaporation of electrons. The corresponding stationary regime in hot dynamical electrosphere can be found by solving the Vlasov-Maxwell system.
@@ III Electrostatic configuration and Schwinger luminosity
which for high degeneracy (when $\eta\gg 1$) reduces to 1995PhRvD..51.1440K ; 2005ApJ…620..915U
@@
Poisson equation for quarks and electrons near the surface gives 1986ApJ…310..261A ; 1995PhRvD..51.1440K
@@
To describe pair creation for non-vacuum initial state it is necessary to take into account Pauli blocking effect 2023PhRvD.108a3002P ; 2023PhRvE.107c5204B . We use the differential pair creation rate given in 1987PhRvD..36..114G with additional Pauli blocking factor $(1-f_{e})$. The rate is then given by an integral over particle momentum $d^{3}p_{e}$ in the following form
@@
Figure 4: Normalized pair rate (16) (blue) with the corresponding normalized electric field (red) for $k_{B}T=m_{e}c^{2}$.
@@
The set of Vlasov-Maxwell equations for distribution functions $f_{\pm}$ and electric field $\tilde{E}$ in our case reduces to the Vlasov-Ampére system 2024ApJ…963..149P
@@
Figure 5: Stationary solution of Vlasov-Maxwell equations for electrosphere for $k_{B}T=2m_{e}c^{2}$ with dependence on distance of electric field (top), particle density (middle) and luminosity (bottom). Approximate electrostatic solution given by eq. ((15)) is shown in black. Colors correspond to: electric field (red), electrons (blue) positrons (orange), and electrostatic solution (black).
@@
IV Heat transfer and neutrino losses
@@
Now we are going to consider heat transfer in the quark star accounting for temperature evolution on much large timescales than the ones studied above. First we treat neutrino transparent case, following 2002PhRvL..89m1101P ; 2001A&A…368..561B ; 2022A&A…663A..19Z ; PhysRevC.85.035805 .
Specific heat capacity $c_{v}$ for unpaired degenerate quarks is IWAMOTO19821
@@
Heat conductivity $\kappa$ for unpaired degenerate quarks is Haensel:1989ja ; 1991NuPhS..24…23H
$\kappa=\frac{3}{\pi^{3}}\sqrt{\frac{3}{2\pi}}\frac{\hbar k_{F}^{3}c^{2}}{\alpha_{s}^{1/2}T},$ (21)
where $k_{F}$ is Fermi wavenumber of quarks, $\alpha_{s}=0.3$ is QCD coupling constant.
In unpaired quark matter neutrino emissivity is dominated by the direct URCA processes of light quarks $d\rightarrow u+e+\bar{\nu}_{e}$ and $u+e\rightarrow d+\nu_{e}$. The energy emission per unit time and volume is IWAMOTO19821
@@
IV.1 Neutrino transparent regime
@@
We solved numerically equation (23) for the stellar configuration with the mass $M=1.4M_{\odot}$. The temperature evolution of the star is shown in figure 6 together with luminosity evolution. After the time moment $t>1$ second, our curves for neutrino demonstrate the same shape as in fig. 1 in 2002PhRvL..89m1101P .
@@
The most important result is that the temperature evolution on the surface is independent on interior of the star. Surface temperature decreases up to few $5\times 10^{9}$ K at $t\simeq 10^{2}$ seconds. The corresponding luminosity is about $10^{43}$ erg/s. The temperature gradient disappears only at $10^{2}$ seconds, when the pair luminosity starts to decrease rapidly.
The total energy emitted in neutrinos is about $10^{53}$ ergs, while the total energy emitted in pairs is only $8\times 10^{46}$ ergs.
@@
IV.2 Neutrino opaque regime
It is believed that neutrinos cannot escape freely immediately after strange star formation, until internal temperature decreases below $10^{10}$ K 2002PhRvL..89m1101P . Following that idea, the integral models for neutrinosphere was developed in 2004csqn.conf..399B ; 2021ApJ…922..214L , where quark star is divided into two parts having different temperatures, separated from each other by a neutrinosphere with radius $R_{\nu}<R$. Neutrinosphere emits blackbody neutrino spectrum from its surface. The results of 2004csqn.conf..399B imply that neutrino opacity delays the process of cooling by an order of magnitude when compared to the transparent case. The results in 2021ApJ…922..214L imply that high luminosity in pairs up to $10^{50}$ erg/s can be sustained for long time, up to $10$ seconds. In this Section we compare the integral model developed in 2021ApJ…922..214L with the differential models adopted in our work.
Figure 7: Surface temperature evolution for neutrino transparent shell of the star with selected positions of neutrinosphere: $R-R_{\nu}=0.001$ cm (red), $R-R_{\nu}=0.01$ cm (orange), $R-R_{\nu}=0.1$ cm (green), $R-R_{\nu}=1$ cm (blue). Black curve represents blue curve from figure 6 (top) for convenience.
@@ IV.2 Neutrino opaque regime
Consider an outer shell of the star and fix the temperature at its inner boundary to be equal to the initial temperature $T(t,r=R_{\nu})=T(t=0,r)=T_{0}$. We solve numerically the heat transfer equation for different thickness of the shell $R-R_{\nu}=\{0.001,0.01,0.1,1\}$ cm. The results for the surface temperature are presented in figure 7. It is clear that the surface temperature decreases very fast in all cases. When the thickness increases up to $R-R_{\nu}=1$ cm cooling occurs in the same way as for neutrino transparent star, that is the surface temperature is about $10^{10}$ K at the time moment $t\simeq 0.2$ seconds. Finally, in figure 8 we show temperature distribution $T(r)$ and heat flux distribution $F_{r}(r)$ (for convenience we plot the value $4\pi R^{2}F_{r}$) inside the shell with thickness about 1 cm at different instants of time. Clearly, initially very steep temperature gradient emerges instantly and remains localized at the surface, due to weak thermal conductivity of quarks.
@@
In realistic model neutrinosphere moves toward the center of the star and at the same time loses its temperature. It means that our model overestimates the surface temperature. Thus, our results imply that the surface temperature of a newborn strange star made of unpaired quarks decreases below the value $10^{10}$ K on the timescales $\Delta t\ll 1$ second, and the reason for that is the very high luminosity due to the Schwinger process and weak thermal conductivity of unpaired quarks.
V Discussion and conclusions